candygirl200:

Trevon can oil lanes in a bowling alley in 7 hours. Imani can oil the same lanes in 10 hours. Find out how long it would take them if they worked together.

Shadow:

So the way that you would first approach this question is to put what Trevor and Imani can do in terms of an hour Trevon = 7hours Trevon can do 1/7 of a lane in an hour Imani = 10hours Imagine can do 1/10 of a lane in an hour \[\frac{ 1 }{ 7 } + \frac{ 1 }{ 10 } = \frac{ 10 }{ 70 } + \frac{ 7 }{ 70 } = \frac{ 17 }{ 70 }\]

Shadow:

Together, they can do 17/70 of a lane, in an hour

Shadow:

Imagine, lol I meant Imani :)

Shadow:

Does this make sense so far?

candygirl200:

yes

Shadow:

Now we need a variable, and an equation, for how long it would take for them to do it together. Let t = time it would take for them to do the job together \[\frac{ 1 }{ t } = \frac{ 17 }{ 70 }\] Flip it \[\frac{ t }{ 1 } = \frac{ 70 }{ 17 }\] \[t = \frac{ 70 }{ 17 }\]

Shadow:

Divide the two numbers, and you should "Imani and Trevor can do the job together in roughly 4.1hours."

Shadow:

Though if your teacher wants a fraction, or a decimal rounded to a certain point, be sure to give that instead.